# PAC-Bayesian Reinforcement Learning Trains Generalizable Policies

### v1

@@ 3.2 Main Result
The bounded-differences property defined above is precisely what allows us to apply concentration inequalities for dependent data. Specifically, we leverage Paulin (2018)’s extension of McDiarmid’s inequality to Markov chains, which provides concentration for functions satisfying bounded differences on Markovian sequences. The key insight is that while the transitions are temporally dependent, the bounded-differences condition with explicit constants $\|c\|^{2}$ enables us to control how perturbations propagate through the dependency structure.
Applying Paulin (2018)’s concentration result yields a tail bound on the deviation $\mathcal{L}(\theta)-\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)$ (see equation 15) that depends explicitly on the mixing time $\tau_{\min}$ of the policy-induced Markov chain. $\tau_{\min}$ is the smallest number of steps after which the distribution of the chain’s state is, in a statistical sense, nearly indistinguishable from its long‑run or stationary distribution in Total Variation distance, no matter where the chain started. In other words, it measures how quickly the chain “forgets” its initial state and becomes well mixed. Combining the tail bound from equation 15 with the standard PAC-Bayesian change-of-measure technique gives our main result:
@@
Let the reward function be bounded in $0,R_{max}]$ and let $\mathcal{M}$ be a (not necessarily time-homogeneous) Markov Decision Process (MDP) induced by any policy $\pi_{\theta}$ such that it satisfies $\tau_{\min}<\infty$. For any prior $\mu$ over $\Pi$, any posterior $\rho$ chosen after interacting with the environment, and any $\delta\in(0,1)$, with probability at least $1-\delta$ over the sample $\mathfrak{D}$ of $T$ trajectories with time horizon $H$:
$\underset{\theta\sim\rho}{\mathbb{E}}\left[\mathcal{L}(\theta)-\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)\right]\;\leq\;\sqrt{\frac{R_{\max}^{2}\,\tau_{\min}\bigl(1-\gamma^{2H}\bigr)}{2T(1-\gamma^{2})}\Bigl(\mathrm{KL}(\rho\|\mu)+\ln\frac{2}{\delta}\Bigr)}.$ (8)
The bound in 3.2 can be straightforwardly converted to a PAC-Bayes lower bound on the true expected value function $\mathbb{E}_{\theta\sim\rho}\left[V_{\pi_{\theta}}\right]$ (see Appendix B, equation 32), using the fact that $\mathcal{L}(\theta)=-V_{\pi_{\theta}}$ (equation 5) and a simple rearrangement of terms. The true expected value is lower-bounded by an empirical estimate minus an uncertainty term that accounts for limited data ($1/T$), temporal correlations ($\tau_{\min}$), and posterior complexity ($\mathrm{KL}(\rho\|\mu)$). This interpretation suggests a natural approach to policy optimization: select the posterior $\rho$ that maximizes this lower bound (equivalently, minimizes the upper bound in equation 8). Such a strategy would automatically balance exploitation (maximizing the empirical value) and theoretically-justified exploration (accounting for uncertainty).
@@ 3.3 Key Improvements and Discussion
Improved Scaling.
Previous PAC-Bayesian bounds for RL suffer from poor scaling with the discount factor. Fard et al. (2012) requires $N>R^{4}/(1-\gamma)^{4}$ total time steps for non-vacuity, while Tasdighi et al. (2025) faces similar constraints. For $\gamma=0.99$, this scaling renders such bounds practically vacuous. Our transition-level analysis achieves substantially better scaling: $T>\frac{R_{\max}^{2}\tau_{\min}(1-\gamma^{2H})}{2(1-\gamma^{2})}$ trajectories suffice. Although this may look prohibitive due to dependence on the number of trajectories, the key improvements are: (i) $\gamma$ appears with power 2 rather than in $(1-\gamma)^{4}$, and (ii) collecting many short trajectories makes the numerator small since $(1-\gamma^{2H})<1$.
@@
The bound requires estimating $\tau_{\min}$, which presents both computational and robustness considerations. We estimate mixing time using autocorrelation decay of the reward signal, as this can be computed from streaming trajectories without storing full visitation counts. Alternative approaches using spectral methods could provide better estimates leading to tighter bounds but require sufficient state-action coverage to construct reliable empirical transition matrices which is infeasible when the state space is continuous.
Our estimation approach is robust to errors in a specific direction: if we overestimate $\tau_{\min}$, our bound remains valid but becomes looser, which is harmless in practice. However, underestimation can be problematic as it leads to overconfidence. To mitigate this, one can compute autocorrelation from multiple sources (e.g., state features, value estimates) to cross-validate mixing time estimates. In practice, we err on the side of caution by using a conservative initial estimate and taking the maximum with the latest autocorrelation estimation, trading some tightness for reliability.
@@ 4 PB-SAC: A Practical Algorithm for PAC-Bayesian Reinforcement Learning
Translating the theoretical PAC-Bayesian bound into a practical learning algorithm requires addressing challenges of maintaining posterior distributions over policy parameters in deep RL. Our algorithm, PAC-Bayes Soft Actor-Critic (PB-SAC), builds upon SAC while integrating PAC-Bayesian bounds and posterior-guided exploration (pseudo-code and illustrative figure in Appendix D).
@@ 4.1 Posterior-Guided Exploration
During exploration, PB-SAC leverages the posterior distribution to implement uncertainty-driven exploration. Rather than standard $\epsilon$-greedy exploration, the algorithm samples policies from the posterior and selects actions that maximize Q-values under posterior uncertainty: $\theta_{explore}\leftarrow\arg\max_{\theta\sim\rho}Q(s,\pi_{\theta}(s))$}. This posterior-guided exploration naturally balances exploitation of the mean policy (the current actor) when it yields the highest value, with exploration of alternative policies in regions of high posterior uncertainty where potentially superior policies may exist. This approach provides theoretical grounding for the exploration strategy through the PAC-Bayesian framework, ensuring that exploration is guided by uncertainty quantification rather than arbitrary randomness.
4.2 Alternating Optimization via PAC-Bayes-$\kappa$
[…]
Our implementation optimizes the following PAC-Bayes-$\kappa$ objective to obtain the posterior $\rho$ then substitutes this optimized posterior into equation 8. As shown in our proof B.5.5, the main bound achieves its minimum over $\kappa^{*}$, yielding a tighter certificate:
$\mathcal{L}(\rho,\kappa)=\underset{\theta\sim\rho}{\mathbb{E}}\left[\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)\right]+\frac{\mathrm{KL}(\rho\|\mu)}{\kappa}+\frac{\kappa\|c\|^{2}\tau_{\min}}{8}$ (9)
This decomposition enables stable alternating optimization: we optimize posterior parameters for fixed $\kappa$, then optimize $\kappa$ for fixed posterior parameters, ensuring convergence while maintaining the theoretical guarantees of our PAC-Bayesian bound.
@@ 4.3 Policy-level REINFORCE trick
Even with the above decomposition, we cannot straightforwardly compute the gradient $\nabla_{(\upsilon,\sigma)}\mathbb{E}_{\theta\sim\rho}\left[\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)\right]$ because sampling cannot occur inside the gradient operation. To address this challenge, we employ a two-stage approach. First, we collect fresh rollouts during the PAC-Bayes update cycle using the mean policy, then sample policies from $\rho$ and evaluate their discounted returns on these rollouts using importance sampling. This ensures our bound computation accounts for distributional shift between the data-generating policy and the current posterior distribution. With estimated returns for each sampled policy, we apply the log-likelihood trick (REINFORCE) (Williams, 1992) at the policy level rather than the traditional action level. We prove this extension in Appendix C. This technique allows us to exchange gradient and expectation operations: $\mathbb{E}_{\theta\sim\rho}\left[\nabla_{(\upsilon,\sigma)}\log\mathbb{P}_{\upsilon,\sigma}(\theta)\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)\right]$, yielding a tractable sampling-based gradient estimator.
4.4 The Adaptive Sampling Curriculum
Our most critical innovation addresses the ”posterior syncing shock” problem: when the posterior mean shifts significantly after PAC-Bayesian updates, critics become misaligned with the new policy distribution, creating a destabilizing feedback loop where inaccurate value estimates generate misleading gradients that further destabilize training through exploding gradients and learning instability.
PB-SAC resolves this through an adaptive sampling curriculum: immediately following PAC-Bayesian updates, we freeze the actor and employ high-rate posterior sampling (512 samples) to expose critics to the full posterior distribution for recalibration, then resume efficient learning with minimal sampling (1 sample, the posterior mean), achieving computational efficiency without sacrificing stability.
@@ 5 Experiments
We now turn to empirical validation across representative continuous control tasks to demonstrate that our PAC-Bayesian framework delivers on this promise while maintaining competitive learning performance.
@@ 5.1 Experimental Design
We evaluate PB-SAC on four MuJoCo continuous control environments (Towers et al., 2024; Tassa et al., 2018) spanning different complexity levels: HalfCheetah, Ant, Hopper, and Walker2d (in Appendix E). Our evaluation tracks two critical metrics: PAC-Bayesian certificate evolution (Figure 1, right) and learning performance relative to baselines (Figure 1, left).
We run PB-SAC with PAC-Bayesian updates every 20,000 steps, employing our adaptive sampling curriculum (512 posterior samples during critic adaptation, posterior mean alone during regular training). We compare against vanilla SAC using identical network architectures and include PBAC Tasdighi et al. (2025), a PAC-Bayes deep exploration method, despite its primary strengths manifesting in sparse reward settings. Table 1 summarizes all hyperparameters.
@@ 5.2 The Tightening of Performance Certificates
Figure 1(b) reveals the most compelling aspect of our results: the PAC-Bayesian bounds consistently tighten throughout training, tracking the improvement in learned policies. In HalfCheetah, the bounds become informative within 100k steps and continue tightening as performance improves, demonstrating that our certificates provide genuine confidence estimates rather than vacuous guarantees, without hindering learning. Hopper exhibits similar bound evolution, though with slower convergence reflecting its increased complexity. Particularly noteworthy is the bounds’ behavior during performance fluctuations: they appropriately widen during periods of high variance while tightening when performance stabilizes (e.g., HalfCheetah between 100k and 200k steps). This stability stems from our decaying moving average prior updates, which prevent KL explosion in early stages; without this mechanism, KL explosion hinders learning by pulling the posterior back toward the prior rather than allowing improvement.
Ant presents the most challenging test case due to its 3D dynamics and large state space. Even here, our bounds remain meaningful throughout training, tightening during improvement periods (within 500k steps) and providing conservative estimates during temporary performance drops (between 800k and 900k steps). The bounds capture the qualitative behavior predicted by our theory: environments with slower mixing yield more gradual certificate improvement.
@@ 5.3 Empirical Validation Across Environments
The learning curves in Figure 1(a) address a fundamental concern about PAC-Bayesian RL: whether theoretical guarantees necessitate performance sacrifices. Our results demonstrate that PB-SAC matches or exceeds SAC’s sample efficiency across all tested environments. In HalfCheetah and Hopper, PB-SAC shows superior performance compared to both SAC and PBAC. This validates our architectural choices, particularly the policy-posterior synchronization that ensures most learning follows proven SAC dynamics. The periodic PAC-Bayesian updates serve as refinements rather than disruptions, guided by the adaptive sampling curriculum that prevents critic destabilization.
In Ant, PBAC initially outperforms both methods thanks to its multi-objective design accounting for diversity, coherence, and propagation, beneficial when rewards are less informative. However, excessive exploration eventually backfires, causing performance degradation in later training. PB-SAC matches its baseline with slightly faster convergence around 500k steps, though lagging slightly at the end. While not designed specifically for deep exploration, it consistently matches baseline performance while providing theoretical guarantees.

### v2

@@ 3.2 Main Result
The above bounded-differences property (Lemma 3.1) is precisely what allows us to apply concentration inequalities to dependent data. Importantly, the proof strategy is leveraging Paulin (2018)’s extension of McDiarmid’s inequality to Markov chains, which provides concentration for functions satisfying bounded differences on Markovian sequences. The key insight is that while the transitions are temporally dependent, the bounded-differences condition with explicit constants $\|c\|^{2}$ enables us to control how perturbations propagate through the dependency structure.
Applying Paulin (2018)’s concentration result yields a tail bound on the deviation $\mathcal{L}(\theta)-\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)$ (see inequality equation 20) that depends explicitly on the mixing time $\tau_{\min}$ of the policy-induced Markov chain. $\tau_{\min}$ is the smallest number of steps after which the distribution of the chain’s state is, in a statistical sense, nearly indistinguishable from its long‑run or stationary distribution in total variation distance, no matter where the chain started. In other words, it measures how quickly the chain “forgets” its initial state and becomes well mixed. Combining the tail bound from inequality equation 20 with the standard PAC-Bayesian change-of-measure technique gives our main result:
@@
Let the weighted reward function be bounded in $0,R_{max}]$ and let $\mathcal{M}$ be a (not necessarily time-homogeneous) Markov Decision Process (MDP) induced by any policy $\pi_{b}$ such that it satisfies $\tau_{\min}<+\infty$. For any prior $\mu$ over $\Theta$, any posterior $\rho$ chosen after interacting with the environment, and any $\delta\in(0,1)$, with probability at least $1-\delta$ over the sample $\mathfrak{D}$ of $T$ trajectories with time horizon $H$, :
$\mathbb{E}_{\theta\sim\rho}[\mathcal{L}(\theta)]$ $\leq\mathbb{E}_{\theta\sim\rho}[\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)]+\sqrt{\frac{R_{\max}^{2}(1-\gamma^{2H})}{T(1-\gamma^{2})}\tau_{\min}\left(\mathrm{KL}(\rho\|\mu)+\ln\frac{\sqrt{2}}{\delta}\right)}.$ (13)
The bound in Theorem 3.2 can be straightforwardly converted to a PAC-Bayes lower bound on the true expected value function $\mathbb{E}_{\theta\sim\rho}\left[V_{\pi_{\theta}}\right]$ (see Appendix B, inequality equation 31), using the fact that $\mathcal{L}(\theta)=-V_{\pi_{\theta}}$ (equation 8, equation 9) and a simple rearrangement of terms. The true expected value is lower-bounded by an empirical estimate minus an uncertainty term that accounts for limited data ($1/T$), temporal correlations ($\tau_{\min}$), and posterior complexity ($\mathrm{KL}(\rho\|\mu)$). This interpretation suggests a natural approach to policy optimization: select the posterior $\rho$ that maximizes this lower bound (equivalently, minimizes the upper bound in inequality equation 13). Such a strategy would automatically balance exploitation (maximizing the empirical value) and theoretically-justified exploration (accounting for uncertainty).
@@ 3.3 Key Improvements and Discussion
Improved Horizon Dependence.
Previous PAC-Bayes bounds for RL suffer from prohibitive scaling with the effective horizon $1/(1-\gamma)$. Fard et al. (2011) bound the value error via the Bellman error, a conversion (their Eq. 7) that degrades sample complexity to $\mathcal{O}((1-\gamma)^{-4})$. Tasdighi et al. (2025) face comparable constraints. For $\gamma=0.99$, such scaling renders these bounds practically vacuous. In contrast, our transition-level analysis bounds the value error directly via the concentration of discounted returns, achieving a significantly tighter scaling of $\mathcal{O}((1-\gamma)^{-1})$. This improvement arises because our sensitivity term scales with $\sum^{H}\gamma^{2h}\approx(1-\gamma)^{-1}$ (Appendix B.2), avoiding the $(1-\gamma)^{-2}$ penalty inherent to the two-step derivation from the Bellman error. Specifically, we require only $T\gtrsim\frac{R_{\max}^{2}\tau_{\min}}{1-\gamma^{2}}$ trajectories.
@@
The bound requires estimating $\tau_{\min}$, which presents both computational and robustness considerations. We estimate mixing time using autocorrelation decay of the reward signal, as this can be computed from streaming trajectories without storing full visitation counts. Alternative approaches utilizing the pseudo-spectral gap Karagulyan & Alquier (2025) can offer tighter bounds when the gap is estimable. However, constructing fully empirical estimators for these gaps currently requires finite state spaces or specific parametric assumptions (e.g., AR(1)), rendering them intractable for general deep reinforcement learning in continuous environments.
Our estimation approach is robust to errors in a specific direction: if we overestimate $\tau_{\min}$, our bound remains valid but becomes looser, which is harmless in practice. However, underestimation can be problematic as it leads to overconfidence. To mitigate this, we compute autocorrelation from additional sources (e.g., state features) to cross-validate mixing time estimates. In practice, we err on the side of caution by using a conservative initial estimate and taking the maximum with the latest autocorrelation estimation, trading some tightness for reliability. Empirical analysis can be found in Section 5.4.
@@ 4 PB-SAC: A Practical Algorithm for PAC-Bayesian Reinforcement Learning
Translating the theoretical PAC-Bayes bound into a practical learning algorithm requires addressing challenges of maintaining posterior distributions over policy parameters in deep RL. The novel algorithm proposed here, PAC-Bayes Soft Actor-Critic (PB-SAC), builds upon SAC while integrating PAC-Bayesian bounds and posterior-guided exploration (pseudo-code and illustrative figure in Appendix F).
@@ 4.1 Posterior-Guided Exploration
During exploration, PB-SAC leverages the posterior distribution to implement uncertainty-driven exploration. Rather than standard $\epsilon$-greedy exploration, the algorithm samples policies from the posterior and selects actions that maximize Q-values under posterior uncertainty:
$\theta_{explore}\leftarrow\arg\max_{\theta_{i}\in\mathfrak{T}}Q(s,\pi_{\theta}(s))$
such that $\mathfrak{T}$ is a set of policy parameter vectors drawn from $\rho$. This posterior-guided exploration naturally balances exploitation of the mean policy (the current actor) when it yields the highest value, with exploration of alternative policies in regions of high posterior uncertainty where potentially superior policies may exist. This approach provides theoretical grounding for the exploration strategy through PAC-Bayesian computations, ensuring that exploration is guided by uncertainty quantification rather than arbitrary randomness.
4.2 Alternating Optimization via PAC-Bayes-$\lambda$
The core challenge in optimizing our PAC-Bayesian bound (Theorem 3.2) lies in its structure: while the KL divergence term $\mathrm{KL}(\rho\|\mu)$ is convex in its first argument (the posterior parameters), the square root composition in the bound is not guaranteed to be convex, potentially leading to optimization difficulties. To address this challenge, we use variational approximation. By applying the identity $\sqrt{x}=\inf_{\lambda>0}(\frac{x}{2\lambda}+\frac{\lambda}{2})$, we recover the convex objective $\mathcal{J}(\rho,\lambda)$ in equation 14. This objective has a structure remarkably similar to the PAC-Bayes-$\lambda$ bound of Thiemann et al. (2017), derived from the classical PAC-Bayes-kl bound for majority vote learning (see, e.g., Germain et al. (2015), Corollary 21; Seeger (2003); Langford (2005)). Both approaches share the key insight of introducing an auxiliary trade-off parameter $\lambda$ that transforms non-convex objectives into quasi-convex relaxations amenable to stable alternating optimization.
Our implementation optimizes the following PAC-Bayes-$\lambda$ objective to obtain the posterior $\rho$ then substitutes this optimized posterior into inequality equation 13. As shown in our proof B.5.4, the main bound achieves its minimum over $\lambda^{*}$, yielding a tighter certificate:
$\mathcal{J}(\rho,\lambda)=\mathbb{E}_{\theta\sim\rho}[\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)]+\frac{\|c\|^{2}\tau_{\min}\left(\mathrm{KL}(\rho\|\mu)+\ln\frac{\sqrt{2}}{\delta}\right)}{2\lambda}+\frac{\lambda}{2}$ (14)
This decomposition enables stable alternating optimization: we optimize posterior parameters for fixed $\lambda$, then compute optimal closed form $\lambda$ for fixed posterior parameters (equation equation 35), ensuring convergence and maintaining the theoretical guarantees of our PAC-Bayesian bound.
@@ 4.3 Policy-level REINFORCE trick
Even with the above decomposition, we cannot straightforwardly compute the gradient
$\nabla_{(\upsilon,\sigma)}\mathbb{E}_{\theta\sim\rho}\left[\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)\right]$
because sampling cannot occur inside the gradient operation. To address this challenge, we employ a two-stage approach. First, we collect fresh rollouts during the PAC-Bayes update cycle using the mean policy, then sample policies from $\rho$ and evaluate their discounted returns on these rollouts using importance sampling. This ensures that bound computation accounts for distributional shift between the data-generating policy and the current posterior distribution. With estimated returns for each sampled policy, we apply the log-likelihood trick (REINFORCE, Williams (1992)) at the policy level rather than the traditional action level. We prove this extension in Appendix E. This technique allows us to exchange gradient and expectation operations:
$\mathbb{E}_{\theta\sim\rho}\left[\nabla_{(\upsilon,\sigma)}\log\mathbb{P}_{\upsilon,\sigma}(\theta)\hat{\mathcal{L}}_{\mathfrak{D}}(\theta)\right],$
yielding a tractable sampling-based gradient estimator.
4.4 Adaptive Sampling
Our most critical innovation addresses the actor-critic misalignment problem that arises after PAC-Bayesian updates. When the posterior mean shifts significantly, critics become misaligned with the new policy distribution, creating instability in training. Without this mechanism, ablation studies (Figure 7(a)) reveal a characteristic sawtooth pattern: performance drops sharply immediately after each PAC-Bayesian update, then gradually recovers until the next update cycle.
PB-SAC resolves this through adaptive sampling: immediately following PAC-Bayesian updates, we freeze the actor and employ high-rate posterior sampling (256 samples) to expose critics to the full posterior distribution for recalibration, then resume efficient learning with minimal sampling (1 sample, the posterior mean), achieving computational efficiency without sacrificing stability.
@@ 5 Experiments
We now turn to empirical validation across representative continuous control tasks to demonstrate that PAC-Bayesian computations for RL deliver on this promise while maintaining competitive learning performance.
@@ 5.1 Experimental Design
We evaluate PB-SAC on four MuJoCo continuous control environments (Towers et al., 2024; Tassa et al., 2018) spanning different complexity levels: HalfCheetah, Ant, Hopper, and Walker2d. Our evaluation tracks two critical metrics: PAC-Bayesian certificate evolution (Figure 1, right) and learning performance relative to baselines (Figure 1, left).
We run PB-SAC with PAC-Bayesian updates every 20,000 steps, employing our adaptive sampling (512 posterior samples during critic adaptation, posterior mean alone during regular training). We compare against vanilla SAC using identical network architectures and include PBAC Tasdighi et al. (2025), a PAC-Bayes deep exploration method, despite its primary strengths manifesting in sparse reward settings. We justify this comparison by evaluating our algorithm in the same sparse-reward setting, with results reported in Appendix G. Additional ablation studies examining the impact of our algorithmic components (adaptive sampling, posterior-guided exploration, and mixing time estimation) are also provided in Appendix G. Table 1 summarizes all hyperparameters.
@@ 5.2 The Tightening of Performance Certificates
Figure 1(b) reveals the most compelling aspect of our results: the PAC-Bayesian bounds consistently tighten throughout training, tracking the improvement in learned policies. In HalfCheetah, the bounds become informative within 100k steps and continue tightening as performance improves, demonstrating that our certificates provide genuine confidence estimates for high-performing policies rather than vacuous guarantees. Particularly noteworthy is the bounds’ behavior during performance fluctuations: they appropriately widen during periods of high variance whilst tightening when performance stabilizes (e.g., HalfCheetah between 100k and 200k steps). This stability stems from our decaying moving average prior updates, which prevent KL explosion in early stages; without this mechanism, KL explosion hinders learning by pulling the posterior back towards the prior rather than allowing improvement.
Hopper, Walker2d (Figure 4), and Ant exhibit similar bound evolution, with certificates tracking performance and reflecting meaningful task completion. Notably, Ant—the most challenging test case due to its 3D dynamics and large state space—demonstrates that our bounds remain meaningful even for complex, high-dimensional continuous control tasks. The bounds capture the qualitative behavior predicted by our theory.
@@ 5.3 Empirical Validation Across Environments
The learning curves in Figure 1(a) address a fundamental concern about PAC-Bayesian RL: whether theoretical guarantees necessitate performance sacrifices. Our results demonstrate that PB-SAC maintains competitive performance with SAC across all tested environments while providing formal certificates. In HalfCheetah and Ant, PB-SAC matches or slightly exceeds SAC performance. In Hopper and Walker2d, both algorithms achieve strong performance, with PB-SAC maintaining stability throughout training. This validates our architectural choices, particularly the policy-posterior synchronization that ensures most learning follows proven SAC dynamics, with periodic PAC-Bayesian updates serving as refinements rather than disruptions, guided by the adaptive sampling that prevents critic destabilization.
PBAC, designed specifically for deep exploration through multi-objective optimization (diversity, coherence, and propagation), underperforms both methods on dense-reward tasks. To ensure a fair comparison—since PBAC’s primary strengths manifest in sparse-reward settings—we evaluated PB-SAC on the sparse-reward task Ant (very delayed) from PBAC’s original work (Tasdighi et al., 2025). Results (Appendix G.1) demonstrate that PB-SAC matches and ultimately surpasses PBAC while outperforming vanilla SAC, confirming that our posterior-guided exploration provides meaningful benefits in sparse-reward settings without requiring PBAC’s high computational complexity (ensemble of critic networks). These results establish that our framework provides non-vacuous certificates and competitive performance across both dense and sparse reward tasks.
5.4 Mixing Time Analysis
A critical component of our framework is mixing time estimation, as underestimation of $\tau_{\min}$ can lead to overconfident bounds. To assess robustness, we conducted experiments with fixed mixing time estimates ranging from 1 (fastest mixing, potential underestimation) to 1000 (conservative overestimation for MuJoCo tasks) without dynamic updates during training. Figure 6 illustrates these results on HalfCheetah. As expected, smaller mixing time values produce tighter bounds [^2] Using a mixing time equal to 1 is equivalent to an i.i.d. bound, more specifically to one that can be derived from the original McDiarmid’s inequality (McDiarmid, 1989) (not the one we used). This demonstrates how tight such bounds can be., while larger values yield more conservative certificates. However, a notable finding is that overestimation proves less problematic than anticipated: while the bound remains theoretically valid but looser, the algorithm effectively optimizes it, maintaining meaningful certificates even with substantial overestimation. This robustness suggests that erring on the side of caution with conservative mixing time estimates is a viable practical strategy. In practice, we mitigate underestimation risk through cross-validation using autocorrelation from multiple signals (rewards and state features) and maintain monotonically increasing estimates by taking the maximum across updates. These results demonstrate that our framework provides reliable certificates across a wide range of mixing time specifications.